OpenStudy (anonymous):

derivative of 3x^2+4x+5/sqrtx

OpenStudy (anonymous):

quotient rule for this one \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=3x^2+4x+5,f'(x)=6x+4, g(x)=\sqrt{x},g'(x)=\frac{1}{2\sqrt{x}}\] algebra will be involved at the end

OpenStudy (anonymous):

ooh, unless it is not a fraction, except at the end is it \[f(x)=3x^2+4x+\frac{5}{\sqrt{x}}\]?

OpenStudy (raffle_snaffle):